Charge needed within a unit sphere centуred at the origin for producing a potential field V=-6r^5/E0 where r less than or equal to 1


1
Expert's answer
2021-01-06T13:53:10-0500

Solution

We know relationship between electric field and potential is given by

"E=-\frac{dV}{dr}=\frac{30r^4}{\epsilon_0}=\frac{q}{4\pi \epsilon_0 r^2}"

Using r=1meter

"q=12pC"